A body of mass 20 kg moving with a velocity of 10 m/s , is suddenly hit by another body moving in the same direction. If the velocity of the former is changed to 15 m/s, what is the impulse by the first body
1
Expert's answer
2020-06-29T14:06:52-0400
"I=\Delta p=m(v_2-v_1)"

"I=20(15-10)=100\ Ns"